# An Efficient Sparse Kernel Generator for O(3)-Equivariant Deep Networks

### v3

@@ 4.4 Acceleration of MACE
Figure 9 (left) compares the rate of molecular dynamics simulation among the different kernel providers. We benchmarked cuE with the optimal data layout for its irreps and included optimizations for symmetric tensor contraction, linear combination layers, and graph convolution. In FP32 precision, we provide 5.3x speeedup over e3nn and 1.5x over cuEquivariance. In FP64 precision, we offer 6.6x speedup over e3nn and 1.43x over cuEquivariance. Because we do not materialize the intermediate tensor product outputs for each edge, our fused kernels reduce the model memory footprint as well.
Figure 9 (right) breaks down the device runtime spent in various kernels traced by the PyTorch profiler. The sum of kernel runtimes in the right figure differs from the times measured in the left figure due to device idle time. Notice that our speedup comes entirely from accelerating the equivariant graph convolution; the CTP kernel and scatter-sum runtime is 4x faster in our implementation compared to cuE in FP32 precision.
By contrast, cuEquivariance also accelerates a symmetric contraction kernel specific to MACE (reducing the runtime counted in “Other” category), and we leave this kernel to future work. Our runtime in the “Other” category can be reduced through torch.compile(), but we did not include this optimization in fairness to cuE (compilation erros in MACE) and e3nn (memory constraints).
Figure 9: Simulation speed of MACE for varying kernel provider and device time breakdown (no torch.compile()).
@@ 5 Conclusions and Further Work
We have established that our sparse kernels achieve consistently high performance on the most common primitives used in equivariant deep neural networks. Our speedup over e3nn and cuEquivariance prove that a sparsity-aware, JIT-compiled kernel can outperform strategies based entirely on dense linear algebra. We see several avenues for future progress:
  • Low Precision Data and MMA Support: Our kernels rely on the single-instruction multiple thread (SIMT) cores for FP32 and FP64 floating point arithmetic. Modern GPUs offer specialized hardware for lower-precision calculation, both using SIMT cores and within matrix-multiply-accumulate (MMA) units. We plan to harness these capabilities in the future.
  • Accelerator and language support: In contrast to cuEquivariance, all of our code is open-source. We hope to cross compile our kernels to HIP for AMD GPUs, as well as offer bindings for JAX and the Julia scientific computing language.
  • Integration into new models: Our software remains accessible to newcomers while delivering the high performance required for massive inference workloads. In conjunction with domain experts, we hope to apply our library to train larger, more expressive equivariant deep neural networks.

### v4

@@ 4.6 Acceleration of Nequip and MACE
Figure 10 (left) compares the rate of molecular dynamics simulation among the different kernel providers. We benchmarked cuE with the optimal data layout for its irreps and included optimizations for symmetric tensor contraction, linear combination layers, and graph convolution. In FP32 precision, we provide a 5.1x speeedup over e3nn and 1.5x over the older v0.2.0 cuE package, noting that the latter does not provide kernel fusion. A similar speedup exists for the FP64 models. Our implementation, however, achieves 0.73x speedup compared to the latest version of cuE that introduces kernel fusion.
To further investigate these benchmarks, Figure 10 (right) breaks down device runtime spent in various kernels. Our time spent on the tensor product (CTP kernels) falls within 2-3 milliseconds of cuE, a highly competitive result. Our performance suffers due to the remaining model components, which contribute to less than 15% of the unoptimized model runtime. To address this, we created a hybrid model (ours-cuE-hybrid) that combines our fused convolution with the linear / symmetric contraction layers offered by cuE. While the hybrid model closes the gap further, the runtime of the remaining kernels is still higher than cuE. This is because the hybrid model preserves the original data layout of MACE layers, whereas cuE transposes several key weight matrices to achieve higher performance. As a consequence, cuE requires a data reordering function for models trained without the package, whereas our kernels have no such restriction.
Figure 10: Simulation speed of MACE for varying kernel provider (left) and device time breakdown (right). cuE-old refers to version 0.2.0 of their package, while the hybrid implementation combines our fused convolution with other model primitives optimized by cuE.
@@ 5 Conclusions and Further Work
We have established that our sparse kernels achieve consistently high performance on several common primitives used in O(3)-equivariant deep neural networks. We see several avenues for future progress:
  • Low Precision Data and MMA Support: Our kernels rely on the single-instruction multiple thread (SIMT) cores for FP32 and FP64 floating point arithmetic. Modern GPUs offer specialized hardware for lower-precision calculation, both using SIMT cores and within matrix-multiply-accumulate (MMA) units. We hope to harness these capabilities in the future.
  • Stable Summation During Convolution: Our kernel generator allows us to easily extend our methods to use stable (Kahan) summation [15] within fused graph convolution. Kahan summation reduces numerical roundoff error during feature vector aggregation across node neighborhoods, promoting energy conservation in simulations.
  • Integration into new models: Our open-source software remains accessible to newcomers while delivering the high performance required for massive workloads. In conjunction with domain experts, we hope to apply our library to train larger, more expressive equivariant deep neural networks.
Acknowledgements and Disclaimers
We thank the anonymous referees for feedback that strengthened our draft. This research was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0022158. This research is also supported by the Office of Science of the DOE under Award Number DE-AC02-05CH11231. We used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award ASCR-ERCAP-33069. We also used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.